# Ginkeepaspectratio=true,clip=true,draft=false

### v1

@@ 4.1.3 Data regions and copiers
The actual choice comes down to performance (incl. a trade-off between the cost of copying and the cost of mapping operations) and the system’s security policy. The advantage of the model is that the decision can be made without affecting the implementation/operation of either the driver or the clients. For example, option (2) is suitable if there is no concern about one client seeing another’s input data (e.g. when all network traffic is encrypted). The Copy component can be inserted transparently: the difference between case (2) and (3) above does not affect the implementation of either the Virt or the client/copier.
@@ 4.1.4 Broadcast
We offer two schemes for handling address resolution request (ARP) packets coming from the network.
The first approach uses a separate ARP client, whose only job is to respond to those requests. This requires that incoming traffic is routed to clients based on MAC address, and that the ARP client has (R/O) access to the MAC-address allocation table.
The alternative approach handles broadcast packets in the RxVirt by enqueuing a copy of the ARP packet in each client’s queue, and reference counting the driver’s buffer containing the packet. We return the buffer to the driver’s RxF queue once each client has enqueued its copy in its RxF queue (and thus the reference count has dropped to zero). Each client is then responsible for handling any broadcast traffic it receives.
@@ 4.1.5 Other device classes
Others, especially storage devices, do not have the clear separation of Tx and Rx traffic of Ethernet, and instead, the storage device only reacts to explicit read or write requests. This results in a slightly different design:
@@
In addition to read and write requests, the Rq may also contain barrier requests, across which the device is not allowed to reorder other requests. Other protocol details support batching of requests. The storage Virt statically partitions the devices between its clients.
The driver exports an information page of device properties, which is appropriately virtualised by the Virt.
@@ 4.2 OS services
For best performance, LionsOS presents a native API that is asynchronous and modelled largely on the device interfaces. For developer convenience and to ease porting of legacy applications, LionsOS also provides a blocking, POSIX-like API that is layered over the native one.
@@
The same approach can be used for storage, by providing a per-client file-system library that directly operates on the virtual storage device provided by the Virt (with an optional copier in between). Alternatively, a single (trusted) file system can be used for all clients, which then is the sole client of the storage Virt.
Sharing across per-client file systems can be enabled by an explicit multiplexer component that connects to multiple clients.
A unified view of multiple storage devices can be provided to a client through a virtual file system that interfaces to multiple per-device file systems. However, multiple storage devices are rare in the embedded space.
Figure 2: File system configuration options.
Figure 2 shows these options. The main take-away is that separation of concerns means that such choices come down to selecting the appropriate components from the “Lego” set and combining them as needed.
@@ 5.1 The starting point: seL4 Microkit
PDs are structured as event handlers, signalling their semaphore will execute the notified function, identifying the sender PD. A server has another handler function, protected, to receive PPCs. Each PD also has an init handler for initialisation.
The system architecture of PDs and their communication channels (semaphores and shared memory regions) is defined in a system description file (SDF). It specifies the ELF files to be loaded into each PD and a PD’s meta-data, including scheduling parameters, access rights to memory regions and caching attributes, and access to interrupts (which appear as semaphores). A PD can monitor a virtual machine, in which case it acts as a private virtual-machine monitor (VMM) which handles virtualisation events from the VM.
@@ 5.2 Queues and state
The design using explicit virtualisers (for separation of concerns) enables another important simplification: All queues are single producer, single consumer (SPSC), lending themselves to simple, lock-free implementations. Specifically, the TxF and RxA queues hold data that is provided by the driver (original producer) and consumed by the client (ultimate consumer), packets flow from right to left in Figure 1; for the TxA and RxF queues the flow is in the opposite direction.
The queues are also inherently bounded, leading to a simple, array-based implementation, where references to particular queue entries are simply array indices. We also require all queues to be a power of two in size, further simplifying implementation and sanitation.
@@ 5.3 Location transparency
In standard producer-consumer fashion, the lock-free SPSC queues are synchronised by semaphores (signalling a Microkit channel). A producer component signals the consumer if new buffers have been enqueued, and the consumer has indicated that it requires signalling by setting a flag that resides in shared memory. A consumer chooses whether or not to set this flag based on if there is meaningful work that it can do with new buffers. In most of our sub-systems it is more common for a producer to be signalling a consumer, but a consumer can also signal a producer if it has done meaningful work, and if the producer has indicated that it needs signalling by setting its shared flag. This can be useful when a producer wishes to enqueue more data in a queue which is currently full.
This approach is completely location transparent: A particular component is not aware whether the component with which it shares a queue is running on the same or a different core. This location transparency of components makes up for the strictly sequential nature of Microkit components: Instead of requiring error-prone, multi-threaded implementation of components to make use of multicore hardware, LionsOS utilises multicore processors by distributing components across cores.
@@
Location transparency will also simplify core management: If a core needs to be off-lined, components running on it can be transparently migrated to other cores, without affecting the system’s operation (other than some temporary latency increases). However, we have not yet implemented this.
@@ 5.4 Legacy driver reuse
The LionsOS design vastly simplifies drivers compared to other OSes (see Section 6.2), and implementing drivers from scratch will result in the best performance.
However, it is unrealistic to expect adopters to write all drivers from scratch, especially since in practice few devices are performance critical enough to justify such an effort. It is also frequently impractical, as many devices are poorly (or un-)documented. For such cases, LionsOS allows reusing a driver from Linux by encapsulating it in a virtual machine (VM). Unlike the Dom0 driver VM of Xen [Barham et al., 2003] and used in other microkernel systems [Chen et al., 2024], we follow the approach of LeVasseur et al. [2004] and wrap each driver in its own VM.
@@ 5.5.1 Device drivers
Most of our development happens on the HardKernel Odroid-C4 (Amlogic S905X3 SoC) and the Avnet MaaXBoard (i.MX8MQ SoC) platforms, so this is where we have the largest set of native drivers. Specifically, we have native drivers for the following device classes:
@@
These are written in C unless otherwise stated, but the system does not prescribe an implementation language, as demonstrated by the Odroid-C4 SDHC driver being written in Rust.
We have Linux-based driver VMs for the following device classes:
@@ 5.5.2 Services
At present we have full networking functionality, using lwIP [Dunkels, 2001] as a client library. Both the native (synchronous) as well as the blocking API are supported, the latter is layered over the former using a coroutine library. We also have an NFS client (using an open-source NFS library) which uses the blocking API. For address-space separation (Section 4.1.3) we use an optional Copy component in the Rx path.
We have an asynchronous filesystem API that is used to talk to a component running NFS or to a VM. The Linux guest in the VM uses the standard VirtIO block interface (converted to our block interface by the VMM), and can make any filesystem supported by Linux available to other components using the filesystem API.
For security, the VM-based filesystem need not be shared between components. Block device sharing is then by partition and enforced by the block virtualiser. Particular use cases can allow more sharing if desired, by using a component between the filesystem and its clients to check allowable access. The VM-based Linux filesystem is not amenable to any kind of formal verification.
The system is complete enough to run several complete systems in daily use. One of them is a reference design for a point-of-sales terminal. It uses a driver VM for re-using the Linux GPU driver, and either a native Ethernet driver or another driver VM for re-using a Linux driver (to demonstrate multiple driver VMs).
@@
Another deployed system is a web server that hosts the sel4.systems web site.
The web server has the business logic implemented in Python, supported by a port of MicroPython [Developers, ] to LionsOS, Figure 4 shows the architecture.
@@ 5.5.3 Resource management
We have not yet implemented dynamic resource management, such as core off-/on-lining (cf. Section 5.3), but plan to do this within a year.
@@ 6.1 Platforms
The Arm platform is an Avnet MaaXBoard with an NXP i.MX8MQ SoC, having four Cortex A53 cores running at a maximum of 1.5 GHz; we run our measurements at a clock rate of 1 GHz. It has 2 GiB of RAM, an on-chip 1 Gb/s NIC, and an on-chip SDHC controller. We perform Linux measurements on this board with a small Buildroot system using kernel version 6.1.0.
The x86 platform is an Intel Xeon® W-1250 six-core CPU running at 3.3 GHz, with an Intel IXGBE X550 10 Gb/s copper NIC. We disable hyperthreading and turbo-boost. For Linux measurements we use a Debian Bullseye userspace and use the “performance” CPU frequency controller to keep the CPU frequency at its maximum value.
@@ 6.2 Complexity and development effort
Table 1: Code sizes in lines-of-code of LionsOS drivers compared to Linux.
@@ 6.2.1 Code size
Our subjective experience is that the LionsOS model dramatically simplifies development of core OS components. A striking example is the i.MX8 network driver, which was the first device driver written to the LionsOS driver model of Section 4.1.1. It was implemented by a second-year undergraduate student, less than 18 months after she wrote her first program. Table 1 compares its size with the Linux driver for the same device. It also compares drivers we use on our x86 platform. We use sloccount [Wheeler, 2001] for all code-size measurements.
@@
Our experience with other LionsOS components is similar, components are small and simple. Table 2 gives a breakdown of the point-of-sales terminal we developed as a demonstrator (see Section 5.5.2). LionsOS, as configured for this application, consists of about 3.1 kSLOC of trusted code, plus 66 kSLOC of untrusted library code that cannot break isolation. All the trusted code is written from scratch.
@@ 6.2.2 Signalling protocols
A potential refinement is to signal only if there is a substantial change to a queue: The producer signals only if the queue becomes non-empty, the consumer only if the queue becomes non-full. Implementing such a protocol correctly however can prove to be difficult, as checking the fullness status of a queue modified by two components in shared memory can introduce race conditions. The components can obtain disagreeing measurements, causing both unnecessary signals or failures to signal at critical times. As well as this, the design can still result in over-signalling as a component’s ability to make progress is usually not dependent on only one queue.
As an example, consider the Ethernet driver. It removes active buffers from the device Rx ring and immediately inserts them into the Virt’s RxA queue. It also removes buffers from the Virt’s RxF queue and inserts them into the hardware ring. This means that the Rx code of the driver can only make progress if one of two conditions are satisfied:
  1. 1.
    The hardware Rx ring is non-empty and has active buffers and the RxVirt’s RxA queue is non-full
  2. 2.
    The Virt’s RxF queue is non-empty and the hardware Rx ring is non-full.
This means that if the Virt signals the driver whenever the RxA queue becomes non-full or the RxF queue becomes non-empty, it may wake the driver only for the latter to find that the hardware Rx ring is not in the necessary state to make progress and go back to sleep.
We address this by letting the driver indicate with a flag for each queue whether it wants to be woken. This allows the driver to incorporate its knowledge of the hardware ring state into the Virt’s decision of whether to signal. However, the resulting increase of complexity of the intertwined protocols is a source of subtle concurrency bugs that can lead to deadlocks.
Fortunately we find that model-checking is a useful tool to alleviate these protocol bugs, and we are able to create models of each of our components using SPIN [Holzmann, 1997], and use these models to prove deadlock freedom of our protocols. This allows us to optimise each component’s decision to signal more aggressively, as we can be assured that no critical signals are missed.
@@ 6.2.3 Driver VMs
Contrary to claims made by Chen et al. [2024] (which are based on work by LeVasseur et al. [2004] that preceded hardware virtualisation support), we find that reusing a Linux driver in a virtual machine is an easy way to gain access to a device. The main challenges are platform initialisation, and avoiding excessive resource consumption at runtime.
@@
The resulting engineering effort per device-class is small (a few weeks). The main disadvantage (apart from the relatively large size of an entire VM for a single driver) is increased latency in accessing the device.
Limiting resource usage
@@
GPU 33 MiB 6.4 MiB 128 MiB
Audio 3 MiB 2.4 MiB 18 MiB
Block 3 MiB 48 KiB 12 MiB
Table 3: Sizes of driver VMs (GPU is not optimised).
RAM usage can be limited by carefully configuring kernel and userspace to use only what is needed (see Table 3). The size differences between the unoptimised GPU and the optimised audio and block device VMs show the order-of-magnitude gains possible with a bit of effort. The resulting memory footprints are small given today’s memory sizes. However, complex SoCs may contain hundreds of devices, so when using a separate VM for each driver, the space cost will add up. We can amortise this cost over multiple devices by supporting multiple drivers from a single VM, at the expense of reduced isolation between drivers – likely a sensible approach for the myriad of simple devices on a complex chip.
@@
Most devices need various platform registers to be set appropriately, to set up clocks for the device, and connections to the outside world (pins). We use native PinMux and clock drivers to set up the platform to enable all devices in the device tree. The driver VM traps accesses to the PinMux and clock registers, discards writes, and passes the actual values in the registers to the native drivers, which provide an interface for querying register values. For debugging, the native drivers support printing the values passed to them.
Some devices require changing PinMux or clock settings at run time. For example Odroid C4 uses a pin either as a clock output, or as a GPIO, while upgrading an SDHC card from its initial low-speed to a high-speed state. Supporting such dynamic setting is future work.
@@ 6.3.1 Networking
Figure 5: LionsOS vs Linux performance on i.MX8MQ UDP echo benchmark, single core. Solid lines are throughput, dashed lines CPU utilisation
Figure 6: LionsOS vs Linux performance on i.MX8MQ UDP echo benchmark, SMP.
@@
To measure networking performance, we configure the evaluation system (LionsOS or Linux) with an echo client, which receives packets from the network and immediately sends them back. The evaluation system also measures CPU load. For LionsOS we use the lwIP protocol stack as a library, linked against the client, and an RxCopy.
@@
We use an external load-generator machine, connected to the evaluation platform via a switch. The load generator runs ipbench [Wienand and Macpherson, 2004] to apply a varying load of UDP traffic and measures the achieved throughput, i.e. the bandwidth received back from the evaluation system. ipbench also measures round-trip time (RTT).
We compare against Linux Debian 6.6.15-2 (2024-02-04) running the standard in-kernel IP stack.
Figure 5 shows the result on the Arm platform, where everything is pinned to a single core. The LionsOS system has no problem keeping up with the load, up to the capacity of the 1 Gb/s NIC, while Linux plateaus at just under 600 Mb/s throughput. Looking at the CPU load we see the reason: Linux maxes out the core at an applied load of 600 Mb/s, while LionsOS achieves the full Gb/s throughput with about 10% CPU capacity left spare. In other words, LionsOS uses just over half the CPU of Linux for handling the same load.
We also run this benchmark on Genode 24.11 [Feske, 2015], using the ”hw” base platform, which runs bare metal without a microkernel underneath. We configure Genode with three components: the echo server (using lwIP), the uplink component (similar to our Virts), and the NIC driver, which is transplanted from Linux 6.6.47. We take all directly from the Genode source tree, compose into a single system with a custom run script based on the mnt_pocket_stmmac_nic example.
The Genode throughput is the blue line in Figure 5. It reaches a maximum of about 100 Mb/s throughput, where it maxes out the CPU, and then collapses at 300 Mb/s applied load. Given this disappointing result, we also run netperf benchmarks on the netperf_lwip example provided by Genode. The TCP_STREAM and TCP_MAERTS benchmarks, which measure unidirectional TCP performance, achieve around 250Mb/s and 230Mb/s respectively, still far below throughputs achieved by either Linux or LionsOS.
Figure 6 shows multicore results on Arm. The story is similar to unicore: Linux fails to handle the full load (but only uses 1.4 cores at maximum load) while LionsOS handles the full load with just over 1.0 cores. CPU load is higher than on unicore as cross-core signalling is more expensive than intra-core.
Figure 7: LionsOS vs Linux performance on x86_64 UDP echo benchmark, single core.
We perform the comparison to Linux on the x86 platform as well, Figure 7 shows unicore results. LionsOS handles the traffic up to an applied load of about 6 Gb/s despite almost maxing out the CPU at a load of 4 Gb/s, after which it can no longer keep up. However, throughput still increases, maxing out at about 7 Gb/s. The continued increase in throughput results from natural batching: the driver finds an increasing number of packets to process per IRQ.
In contrast, Linux, while similarly maxing out the CPU at around 4 Gb/s, shows a performance collapse at increasing load, dropping from a maximum throughput of 3.5 Gb/a to less than 2 Gb/s at maximum applied load.
Figure 8: Single core median round-trip times on x86_64 for LionsOS and Linux, bars show observed range. Dashed lines represent the CPU cost per unit of data.
Figure 8 compares round-trip times for the two systems. LionsOS RTT is consistently low (below 200 $\mu$s) until it increases sharply when the load exceeds 6 Gb/s. In contrast, Linux’ RTT is consistently much higher, around 1000 $\mu$s across the range. Looking at the CPU cycles per KiB processed, LionsOS and Linux start out the same but then diverge, as LionsOS per-packet processing cost drops (due to batching) with increasing load, while Linux’ cost increases.
@@
Figure 9: LionsOS vs Linux performance on x86_64 UDP echo benchmark, SMP.
As can be see from Figure 9 LionsOS can achieve full network bandwidth using two cores, whereas Linux achieves only 4Gb/s (but avoids a performance collapse).
Figure 10: Single core throughput and CPU utilisation on x86_64 for seL4-based compared with simulated Fiasco- and Zircon-basedLionsOS.

### v2

@@ 4.1.3 Data regions and copiers
The actual choice comes down to performance (incl. a trade-off between the cost of copying and the cost of mapping operations) and the system’s security policy. For example, option (2) is suitable if there is no concern about one client seeing another’s input data (e.g. when all network traffic is encrypted) and clients can be trusted to return buffers. The Copy component can be inserted transparently: the difference between case (2) and (3) above does not affect the implementation of either the Virt or the client/copier.
@@ 4.1.4 Broadcast
Broadcast packets such as for the Address Resolution Protocol (ARP) need special handling. We offer two schemes.
The first approach uses a separate ARP client, whose only job is to respond to those requests. This requires that incoming traffic is routed to clients based on MAC address, and that the ARP client has (R/O) access to the MAC-address allocation table. Other broadcast packets are dropped.
The alternative approach handles a broadcast packet in the RxVirt by enqueuing a copy of the packet in each client’s queue, and reference counting the driver’s buffer containing the packet. We decrement the reference count when the buffer is returned to the client’s RxF queue; and return the buffer to the driver’s RxF queue once all clients have finished (i.e. the reference count has dropped to zero). Each client is then responsible for handling any broadcast traffic it receives.
@@ 4.1.5 Other device classes
Others, especially storage devices, do not have the clear separation of Tx and Rx traffic of Ethernet, and instead react to explicit requests. This results in a slightly different design:
@@
For storage, in addition to read and write requests, the Rq may also contain barrier requests, across which the device is not allowed to reorder other requests. Other protocol details support batching of requests. The storage Virt statically partitions the devices between its clients.
The storage driver exports an information page of device properties, which is appropriately virtualised by the Virt.
@@ 4.2 OS services
For best performance, LionsOS presents a native API that is asynchronous and modelled largely on the device interfaces. For developer convenience and to ease porting of legacy applications, we use a coroutine library that implements a POSIX-like blocking API that is layered over the native one.
@@
We use the same approach for storage, by providing a per-client file-system library that directly operates on the virtual storage device provided by the Virt (with an optional copier in between). Alternatively, a single file system could be used for all clients, which then is the sole client of the storage Virt – this shared file system would have to be trusted. We currently see no need for this in our target domain.
Sharing across per-client file systems is enabled by an explicit multiplexer component that connects to multiple clients. In our space, this is generally used for read-only storage.
@@ 5.1 The starting point: seL4 Microkit
PDs are structured as event handlers. Signalling a PD’s semaphore will cause it (eventually) to execute the notified function, identifying the sender PD. A server has another handler function, protected, to receive PPCs. Each PD also has an init handler for initialisation.
The system architecture of PDs and their communication channels (semaphores and shared memory regions) is defined in a system description file (SDF). It specifies the ELF files to be loaded into each PD and a PD’s meta-data, including scheduling parameters, access rights to memory regions and caching attributes, and access to interrupts (which appear as semaphores). A PD can monitor a virtual machine, in which case it acts as a private virtual-machine monitor (VMM) which handles virtualisation events from that VM.
@@ 5.2 Queues and state
The design using explicit virtualisers (for separation of concerns) enables another important simplification: All shared-memory communication is single producer, single consumer (SPSC), enabling the use of simple, lock-free queue implementations. Specifically, the TxF and RxA queues hold data that is provided by the driver (original producer) and consumed by the client (ultimate consumer), packets flow from right to left in Figure 1; for the TxA and RxF queues the flow is in the opposite direction.
The queues are also inherently bounded, leading to a simple, array-based implementation, where references to particular queue entries are array indices. We also require all queues to be a power of two in size, further simplifying implementation and sanitation.
@@ 5.3 Location transparency
In standard producer-consumer fashion, the lock-free SPSC queues are synchronised by semaphores (signalling a Microkit channel). A producer component signals the consumer if new buffers have been enqueued in a previously empty queue, and the consumer has set a flag that requests signalling. Similarly, producers can request signalling on a queue becoming non-full.
This approach is completely location transparent: A particular component is not aware whether the component with which it shares a queue is running on the same or a different core. This location transparency of components makes up for the strictly sequential nature of LionsOS components: Instead of requiring error-prone, multi-threaded implementation of components to make use of multicore hardware, LionsOS utilises multicore processors by distributing components across cores.
@@
Location transparency will also simplify core management (the implementation of which is in progress): If a core needs to be off-lined, components running on it can be transparently migrated to other cores, without affecting the system’s operation (other than some temporary latency increases).
@@ 5.4 Legacy driver reuse
The LionsOS design vastly simplifies drivers compared to other OSes (see Section 6.2); implementing drivers from scratch is usually easy and will result in the best performance.
However, it is unrealistic to expect adopters to write all drivers from scratch, especially since in practice few devices are performance critical enough to justify such an effort. It is also frequently impractical, as many devices are poorly (or un-)documented. For such cases, LionsOS allows reusing a driver from Linux by encapsulating it in a virtual machine (VM). Unlike the Dom0 driver VM of Xen [Barham et al., 2003] or the Driver Container of HongMeng [Chen et al., 2024], we follow the approach of LeVasseur et al. [2004] and support (but do not force) wrapping each driver in its own VM.
@@ 5.5.1 Device drivers
Most of our development happens on the HardKernel Odroid-C4 (Amlogic S905X3 SoC) and the Avnet MaaXBoard (i.MX8MQ SoC) platforms, so this is where we currently have the largest set of native drivers:
@@
Most drivers are written in C, but the system does not prescribe an implementation language, as demonstrated by the Odroid-C4 Rust-implemented SDHC driver.
We have Linux driver VMs for the following device classes:
@@ 5.5.2 Services
We have full networking functionality as described above, using lwIP [Dunkels, 2001] as a client library: lwIP cannot break isolation and is thus not part of the system’s TCB. Our default setup uses an RxCopy component (see Section 4.1.3). Both the native (asynchronous) as well as the blocking API are supported, the latter is layered over the former using a coroutine library. We have an NFS client (using an open-source NFS library) which uses the blocking API.
We have an asynchronous filesystem API that uses either a native FAT file system, the NFS client for network storage, or any Linux file system hosted in a Linux VM accessed via the standard VirtIO block interface. Again, there is a blocking API layered on top.
LionsOS is mature enough to run several complete systems in daily use. One of them is a reference design for a point-of-sales terminal. It uses a driver VM to re-use the Linux GPU driver, and either a native Ethernet driver or another driver VM for re-using a Linux driver (to demonstrate multiple driver VMs).
@@
Another deployed system is a web server that hosts the sel4.systems web site. The web server has the business logic implemented in Python, supported by a port of MicroPython [MicroPython Developers, 2014] to LionsOS, Figure 3 shows the architecture.
@@ 5.5.3 Resource management
Implementing dynamic resource management, such as core off-/on-lining (cf. Section 5.3), is in progress.
@@ 6.1 Platforms
The Arm platform is an Avnet MaaXBoard with an NXP i.MX8MQ SoC, having four Cortex A53 cores running at a maximum of 1.5 GHz and sharing an L2 cache. We run all measurements at a clock rate of 1 GHz to prevent overheating. The board has 2 GiB of RAM, an on-chip 1 Gb/s NIC, and an on-chip SDHC controller. We perform Linux measurements on this board with a small Buildroot [2016] system using kernel version 6.1.0. Some measurements are made on a HardKernel Odroid-C4, which has an AMlogic S905X3 quad-core Cortex A55 running at 1.2 GHz, 4Gb of DDR4 RAM, and on-chip NIC and SDHC devices.
The x86 platform is an Intel Xeon® W-1250 six-core CPU running at 3.3 GHz, private L2 caches and a shared L3. It has an Intel IXGBE X550 10 Gb/s copper NIC. We disable hyperthreading and turbo-boost. For Linux measurements we use a Debian Bullseye userspace and use the “performance” CPU frequency controller to keep the CPU frequency at its maximum value. The kernel is the Debian Linux kernel 6.6.15-2 (2024-02-04) running the standard in-kernel IP stack.
@@ 6.2 Complexity and development effort
Table 1: SLOC of LionsOS NIC drivers compared with Linux.
@@ 6.2.1 Code size
Our subjective experience is that the LionsOS model dramatically simplifies development of core OS components. A striking example is the i.MX8 network driver, which was the first device driver written to the LionsOS driver model of Section 4.1.1. It was implemented by a second-year undergraduate student, less than 18 months after she wrote her first program. Table 1 compares the code sizes of Ethernet drivers to Linux. We use sloccount [Wheeler, 2001] for all code-size measurements.
@@
Our experience with other LionsOS components is similar, components are small and simple. Table 2 gives a breakdown of the Odroid-C4-based point-of-sales terminal (see Section 5.5.2). LionsOS, as configured for this application, consists of about 3.1 kSLOC of trusted code, plus 66 kSLOC of untrusted library code that cannot break isolation. All the trusted code is written from scratch.
@@ 6.2.2 Signalling protocols
A potential refinement is to signal only if there is a substantial change to a queue: The producer adding to an empty queue or the consumer removing from a full queue. However, race conditions can result in subtle protocol bugs that lead to deadlocks.
Fortunately we find that model-checking is effective at eliminating these bugs. We create models of each our components and use SPIN [Holzmann, 1997] to prove deadlock freedom. This allows us to optimise each component’s signalling decision.
@@ 6.2.3 Driver VMs
Contrary to claims made by Chen et al. [2024] (which are based on work by LeVasseur et al. [2004] that preceded hardware support for virtualisation), we find that reusing a Linux driver in a virtual machine is an easy way to gain access to a driver. The main challenges are platform initialisation, and avoiding excessive resource usage at runtime.
@@
The resulting engineering effort per device-class is small (a few weeks). The main disadvantage (apart from the relatively large size of an entire VM for a single driver) is increased latency in accessing the device, so this is unlikely to be an option for performance-critical devices. In practice, the cost of interacting with a driver only matters for a small number of devices (mostly networking, storage and busses).
Reducing resource usage
@@
Default config 29 6.7 70
GPU 33 6.4 128
Audio 3 2.4 18
Block 3 0.05 12
Table 3: Sizes of driver VMs in MiB (GPU is not optimised).
RAM usage is reduced by carefully configuring a minimal kernel and userspace, see Table 3. The default configuration is what the buildroot tool provides for Odroid-C4; the unoptimised GPU configuration has extra memory mapped in for the framebuffer device communication, and is built with all drivers compiled in (rather than modules in the file system). The optimised audio and block device VMs show the order-of-magnitude gains possible with a bit of effort. The memory footprints of the optimised VMs are small given today’s memory sizes (but co-locating uncritical drivers in a single VM is always possible).
@@
Most devices need various platform registers to be configured, to set up clocks for the device, and connections to the outside world (pins). We use native PinMux and clock drivers to set up the platform to enable all devices in the device tree. Linux also wants to control these, so the driver VM traps accesses to the PinMux and clock registers, discards writes, and on read, queries the native drivers, and passes the actual values in the registers to Linux drivers. For debugging, the native drivers support printing the values passed to them.
Some devices require changing PinMux or clock settings at run time. For example Odroid-C4 uses a pin either as a clock output, or as a GPIO, while upgrading an SDHC card from its initial low-speed to a high-speed state. Supporting such dynamic setting is on-going.
@@ 6.3 Performance
The fine-grained modularity of the LionsOS design implies a higher context-switching rate than for monolithic designs, making IPC costs a potential performance bottleneck  [Mi et al., 2019; Chen et al., 2024].
We focus on networking, which stresses the design the most out of all the realistic use cases we can think of: With standard 1.5 KiB Ethernet frames, a 1 Gb/s NIC (typical of embedded platforms) can handle 81,000 packets per second for transmit and the same number for receive, resulting in 100,000s context switches per second; a 10 Gb/s NIC will impose ten times that load. This is at least as challenging as the 60,000 syscalls/s quoted by Chen et al. [2024] for smartphones.
We also look at storage, which can be performance-critical too [Chen et al., 2024].
@@ 6.3.1 Networking
(a) Unicore CPU utilisation.
(b) Unicore round-trip times.
(c) Multicore CPU utilisation.
(d) Multicore CEOS.
Figure 4: Performance comparisons on i.MX8MQ UDP echo benchmark. Here and in other figures, the left axis and solid lines are throughput except where stated otherwise, dashed lines as per graph label (right axis).
@@
To measure networking performance, we configure the evaluated system with an echo client, which receives packets from the network and immediately sends them back.
@@
We use an external load-generator machine, connected to the evaluation platform via a switch. The load generator runs ipbench [Wienand and Macpherson, 2004] to apply a varying load of UDP traffic and measures the achieved throughput, i.e. the bandwidth received back from the evaluation system, as well as round-trip time (RTT). We measure the CPU utilisation of the evaluated system by running on each core a low-priority idle thread that counts free cycles.
(a) Unicore CPU utilisation.
(b) Unicore round-trip times.
(c) Multicore CPU utilisation.
(d) Simulated IPC costs.
@@
Figure 5: LionsOS vs Linux performance on x86_64 UDP echo benchmark. Dashed lines are CPU use except in graph b (RTT).
Linux on Arm
4(a) shows the result on the Arm platform, where everything is pinned to a single core (ignore the blue lines for now). The LionsOS system has no problem keeping up with the load, up to the capacity of the 1 Gb/s NIC, while Linux plateaus at about 500 Mb/s throughput. Looking at the CPU load we see the reason: Linux maxes out the core at an applied load of 500 Mb/s, while LionsOS achieves the full Gb/s throughput with about one third of CPU capacity left spare. In other words, LionsOS uses less than half the CPU of Linux for handling the same load.
4(b) shows that this performance does not come at the expense of latency – LionsOS RTTs are between three and ten times less than Linux’s. In particular, we find that (unlike Linux) we do not need to configure a Rx IRQ hold-off value to handle the load. We do benefit from natural batching, where further packets arrive while the driver is processing earlier ones. When it finishes, the driver checks for new pending interrupts before going to sleep. It will never busy-poll (Linux switches to polling under high load).
4(c) shows multicore results. Here we distribute the LionsOS system over four cores: two cores each run a client and its RxCopy (only one client is active in most benchmarks), one runs the NIC driver and TxVirt, and the forth core runs the RxVirt as well as other system components (timer and serial drivers). This forces control for each packet to cross cores four times per round-trip. Linux is allowed to use cores as it chooses.
We see that the LionsOS CPU load increases from 64% to 78%, a result of cross-core IPC costs being significantly higher than intra-core. Linux still fails to handle the full load (but only uses 1.75 cores at maximum load). The performance gap between the systems is even higher than on unicore.
Other OSes
We also run this benchmark on Genode 24.11 [Feske, 2015], the seemingly only other highly modular protected-mode OS. Only the ”hw” base configuration runs on our platform. It is the configuration with least overhead, as it runs bare metal without a microkernel underneath, so some system services run in privileged mode with no IPC overhead, as in a monolithic kernel. Key networking components, such as the IP stack, resource multiplexers and the NIC driver, still run in userspace.
We configure Genode with three components on unicore: the echo server (using lwIP), the uplink component (similar to our Virts), and the NIC driver, which is transplanted from Linux 6.6.47. We take all directly from the Genode source tree, compose into a single system with a custom run script based on the mnt_pocket_stmmac_nic example. lwIP could be easily swapped out for the other provided network stack, lxIP, which is ported from Linux, but we found that this performed significantly worse in our experiments.
The Genode throughput is the blue line in 4(a). It reaches a maximum of about 100 Mb/s throughput, where it maxes out the CPU, and then collapses at 300 Mb/s applied load. Given this disappointing result, we also run netperf benchmarks on the netperf_lwip example provided by Genode. The TCP_STREAM and TCP_MAERTS benchmarks, which measure unidirectional TCP performance, achieve around 250Mb/s and 230Mb/s respectively, still far below throughputs achieved by either Linux or LionsOS. [^1] We asked for help on the Genode mailing list and communicated directly with Genode staff, and followed their suggestions, which did not improve performance above what we report here.
We also compare to a commercial microkernel-based operating system, code-named CEOS. [^2] The licensing conditions do not allow us to identify the system. CEOS does not support our standard platform (the Maaxboard.), so we evaluate it and LionsOS on the Compulab IOTGate (i.MX 8M Plus SoC), with 4 Cortex-A53 cores running at 1.2 GHz and 6 GiB of RAM. The board has two 1 Gb/s NICs; in our evaluation, we use the same FEC NIC that is used by the Maaxboard. The CEOS network driver is provided in binary format and does not use certain hardware features, such as Tx IRQ coalescing and HW checksum offloading, so we disable these features in our LionsOS driver for fair comparison. We run the benchmark on SMP configurations of both CEOS and LionsOS, as the CEOS networking architecture is optimized for multicore and performed poorly in single-core experiments.
CEOS provides a POSIX socket API, which is the suggested method to interact with the networking architecture, so we port our Linux echo server benchmark to it. This achieves a maximum throughput of around 240 Mbps, as shown by the blue line in 4(d). As the echo server benchmark does not fully utilize the CPU of the system, we also run iperf3, shown by the red line. It is configured as a bidirectional UDP benchmark, with CEOS acting as the server. This achieves better throughput as CPU utilization continues to scale linearly, maxing out at around 390 Mbps, but performance is still far below that of LionsOS (green lines).
(a) Optional components.
(b) Multiple clients, 1 loaded.
(c) Multiple clients, all loaded.
(d) Virtualiser swap experiment.
Figure 6: Component performance investigations.
x86
As Linux might be more performance-tuned on x86 than on Arm, we repeat these measurements on the x86 platform with a 10 Gb/s NIC. 5(a) shows unicore results.
LionsOS again handles the load with about 20% CPU to spare, while Linux maxes out the CPU at 4 Gb/s load and above that its achieved throughput deteriorates. LionsOS RTTs are again much less than Linux, as 5(b) shows.
The multicore configuration, 5(c), shows that LionsOS CPU usage is almost doubled compared to single-core, apparently a result of cross-core communication being much more expensive on the x86 platform (with per-core L2 caches) than the Arm platform (shared L2), but the system has no problem handling the load. Linux, in turn, does not achieve better throughput than on unicore (but avoids the performance collapse), despite having plenty of CPU head space. We run Linux with a second configuration, where we split the echo server into two components, with separate MAC addresses, hence using separate NIC packet rings (black lines, labelled “2 threads” in the figure). This almost doubles the throughput achieved by Linux, but still leaves it behind LionsOS.